OpenStudy (anonymous):

What is the 7th number in the pattern: 1, 4, 9, 16, 25, …?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

it counts up from odd numbers starting with 3 please medal me

OpenStudy (zepp):

In fact, the pattern here is \[\huge1^2,2^2,3^2,4^2,5^2...\]

OpenStudy (anonymous):

It appears to be a sequence of the squares of natural numbers: 1^2,2^2,3^2,... Hence the answer would be (5+1)^2=6^2=36. However, the general term of this sequence is not given, so this question is actually pointless.

OpenStudy (zepp):

So the 7th number will be \[\large 7^2,~which~is~49\]

OpenStudy (anonymous):

it could be either one actually

OpenStudy (anonymous):

@zepp is correct here. The question asks for the 7th term, not the next term. The next term is 36, the 7th term is 49.

OpenStudy (anonymous):

oh my bad i thought they said next

OpenStudy (anonymous):

Oops..

OpenStudy (anonymous):

yes indeed its is 49